INAF queue observing, 2018 Jan 24-25

We (the observers) take this space to bring attention of the world to the the awkward situation of large part of research personnel in INAF. We are now performing observations for INAF researchers with LBT, one of the biggest telescope in the world. But we are also valued for our scientific contribution, for results that we publish as first authors and for those in which we are involved.

However, we believe that like for many other INAF post-docs, our contract situation and our expectations do not reflect the great results that we, and INAF thanks to us, are obtaing. Like many other post-docs, we are paid on contracts renewed every year pending the availability of external funds.

Moreover, we are not entitled to officially say a word on institutional choice and in many aspects we are not considered as INAF employees, but like external collaborators. Indeed, we cannot even lead INAF projects and build a independent research.

Now, we believe INAF has given the power and instruments to start changing this situation.

Observer: R. Carini, A. Rossi *2 among 400 post-doc of INAF*
Telescope Operator: G. Bechetti
Support Astronomer: B. Rothberg on call
Instrument: LUCI1+LUCI2

Summary

Night shared with LINC team.

Instruments: LBC works, LUCI1 only imaging AO, LUCI2 imaging, LS and MOS, MODS1 ok both channels, and MODS2 blue only

Night duration (12\xB0 el twilight) 2:00am- 13:00pm UT = 11 h 15m

Night shared with LINC team. INAF time 2:00am- 7:30pm UT = 5 h 30m

Observing time: 5 h 15m (87.5%) we have finished 30 minutes later (at about 8), we will give back this time to LN team in the next days

Weather time loss: * *h (0%)

technical time loss 0.75 h (12.5%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
darks darks LUCI2 -- mer,int 100 x 3, 5 exposures ok for Saracco science. for telluric still to be done
telluric Nisini field1 HIP16652 LUCI2 K, J ,H    
Nisini field 1 LUCI2 J K 450*8=1h H to be done. not all stars in the slits. 2 stars on top are out in all +2" offsets. Note that we corrected aligment manually
telluric Nisini for field 3 HIP17971 LUCI2 K, J,H    
Nisini field 3 LUCI2 K, J first 2 450*6=0.75h to be completed
telluric Piconcelli A4 HIP53735 LUCI2 HK   completed
Piconcelli A4 J1027+3543 LUCI2 HKspec 0.3h completed

Details

01:30 start with LUCI2 only

01:54 slew to telluric for Nisini luci2
  • seeing 0.9 on guider
  • luci2.0004-0006 acquisition
  • luci2.0007-0018 spectra

02:30 slew to field 1 for Nisini luci2
  • preset fail, it is too close to zenith, rotators cannot run so fast. wait a little bit for target to go down 02:36 send again
  • seeing 0.8" at 02:39
  • luci2.0019-00 acquisition
  • 45min* TECHNICAL LOST difficult allign the mask, perhaps field stop misallignement
  • done AFC allignement*
  • 03:26 acuisition again
  • luci2.0027-0033 acquisition
  • problem with mask allignment. needs -1.5 degrees rotation and manual offsets. <\br> Also 1 allignment star is not used by script but it should according to finding chart! * this is because one of the three alignment stars is not even highlighted in the acquisition process. The others are too close for computing the rotation properly!
  • 3:58 finally at start science
  • NOTE We fear that 3 stars on top are not in the slit. Others are perfectly in the center
  • luci2.0034-0041 science spectra
  • seeing is 1.3-2 on guider. Variable and poor seeing cit: Geno
  • 4:12 seeing is 1.8 on guider
  • 4:51 seeing is 1.2-1.5 on guider
  • stop after J band. We did all K band and all J band only.

05:02 slew to Nisini field 3 telluric
  • 5:03 seeing is 1.2 on guider
  • luci2.0042-0044 acquisition
  • luci2.0045-0056 spectra

05:38 slew to Nisini field 3
  • 5:43 preset error, send again
  • 5:44 seeing is 1.1" on guider
  • luci2.0057-0062 acquisition. not perfect allignement. adjusted manually, +0.2 rotation and -0.12 dx
  • 06:10* go spectra
  • luci2.0063-0068 spectra We did all K band and first 2 offsets of J
  • 6:11 seeing is 1.4" on guider

07:00 slew to Piconcelli telluric HIP53735
  • luci2.0069-0072 acquisition
  • luci2.0073-0074 spectra
  • 7:22 seeing is 0.8" on guider

07:14 slew to Piconcelli A4 J1027+3543
  • preset error, check pointing, send again
  • luci2.0075-0078 acquisition
  • luci2.0079-0082 spectra
  • 7:29 seeing is 0.7" on guider

07:50 we finished. Telescope to LN team

-- %USERSIG{AndreaRossi - 2018-01-25}%

Comments

 
Topic revision: r8 - 01 Feb 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback